John is 3 times as old as martha, Seven years ago the sum of their ages was 54. How old are they now
Answer by sachi(548)

John is 3 times as old as martha,
let martha is x yrs
john is 3x yrs
Seven years ago
martha is x-7 yrs
john is 3x-7 yrs
the sum of their ages was 54.
so x-7+3x-7=54
or 4x=54+14=68
or x=17 yrs
and john is 51 yrs